# Leveraging Explicit Reasoning for Inference Integration in Commonsense-Augmented Dialogue Models

### v1

@@ 1 Introduction
As such, the goal of commonsense-augmented dialogue modeling is to leverage useful commonsense inferences for producing more compelling and human-like responses in human-computer dialogue. For a given dialogue situation, there are numerous commonsense inferences that can be drawn since inferential commonsense has a many-to-many mapping due to its speculative nature Shen et al. (2022); Finch and Choi (2024). Consequently, commonsense-augmented dialogue modeling is a reasoning process of (1) deriving commonsense knowledge that likely holds true for a given dialogue context, (2) identifying a subset of true commonsense that is appropriate for generating a response, and (3) synthesizing a response from the identified commonsense knowledge. No previous work tackles all three of these components of commonsense-augmented dialogue modeling in an explicit manner. Often, the identification of appropriate commonsense for response generation is tackled jointly with response generation itself, where a generative model’s attention mechanisms implicitly learn which commonsense knowledge is most relevant to generating a response in each dialogue context Zhang et al. (2020); Sabour et al. (2022); Liu and Kilicoglu (2023). However, recent work in a variety of text generation tasks, including dialogue modeling, suggests that breaking down a model into a sequence of explicit reasoning steps improves the correctness and overall quality of model outputs Wei et al. (2022); Chae et al. (2023); Lee et al. (2023)
@@
We find strong evidence that explicit reasoning over commonsense inferences leads to better dialogue interactions, improving the naturalness, engagingness, specificity, and overall quality of the responses against several alternative strategies for dialogue response generation. Further analysis provides insights on which types of commonsense are most useful for response generation, revealing that predictions of personal characteristics and likely future events lead to the most favorable outcomes. Moreover, we assess human-provided explanations of response quality to identify the key response attributes that shape human perception of response quality and the particular response traits enhanced through commonsense integration. Taken together, our results demonstrate that isolating commonsense reasoning into explicit reasoning steps improves several aspects of response quality, achieving a new state-of-the-art in commonsense-augmented dialogue modeling. We make all code, models, and data publicly available to facilitate future work. [^1] https://github.com/sfillwo/DialCS-Reasoning
@@ 2 Related Work
To explore the impact of explicit reasoning over commonsense inferences against the typically utilized implicit reasoning approach, we develop two prompt-based LLM strategies for response generation that treat generated inferences as speculative thoughts to guide follow-up response generation. These two approaches utilize the same inference generation procedure and differ on the strategy for integrating commonsense inferences into a follow-up response. The implicit reasoning variant provides all commonsense inferences as input to an LLM which is prompted to consider the inferences when generating the best follow-up response, similar to previous works. The explicit reasoning variant involves a three-step generate-select-and-respond procedure, using LLM prompting to explicitly identify the best commonsense inferences and subsequently synthesize them into a follow-up response. We first discuss the shared Inference Generation module (§3.1), before detailing the implicit reasoning (§3.2) and the explicit reasoning approaches (§3.3). Figure 2 presents an overview of the approaches.
@@ 3.1 Inference Generation
10 commonsense inferences are outputted for a provided dialogue context, each corresponding to one of the 10 commonsense types covered by ConvoSenseGenerator (Table 1). Initially, we explored outputting the top-ranked inference from beam search for each type but found significant semantic overlap in the inferences outputted across types. Since semantically unique inferences are critical for studying the impact of reasoning over these inferences, we implement a diverse beam search approach Vijayakumar et al. (2016) to output five inferences per type and then select one inference per type such that between-type cosine similarity of SBERT inference embeddings Reimers and Gurevych (2019) is minimized, following the results in Finch and Choi (2024) that show diverse beam search helps to increase inference uniqueness.
@@ 3.2 Implicit Reasoning
Given the output from inference generation, the implicit reasoning approach immediately performs response generation by taking all generated inferences as input. Table 6 provides the prompt for this approach (Appx. A), referred to as ConvoSense-I.
@@ 3.3 Explicit Reasoning
Given the output from inference generation, the explicit reasoning approach performs inference selection over the generated inferences before composing the follow-up response by expressing the selected inferences. Table 5 provides the prompts for this approach (Appx. A), referred to as ConvoSense-E.
@@ 3.4 Prompt Formatting
For all GPT prompts, the dialogue context is provided as a sequence of turns, prefixed with speaker labels. Few-shot examples are also provided to both ConvoSense-E and ConvoSense-I. For ConvoSense-E, we construct 10 inference selection examples, one for each inference type included in this study, and 100 response generation examples, 10 for each inference type. During inference, the 10 inference selection examples are used in the inference selection prompt. Once an inference is selected from inference selection, then the response generation examples corresponding to its type are used in the response generation prompt (Table 5; Appx. A). For ConvoSense-I, we choose 10 response generation examples from those crafted for ConvoSense-E, ensuring one example for each commonsense type. During inference, these 10 response generation examples are used in the response generation prompt (Table 6; Appx. A).
@@ 4 Experiments
To study the impact of explicit reasoning over commonsense inferences on response generation, we compare ConvoSense-E against three alternative approaches, two which utilize implicit reasoning over commonsense (ConvoSense-I and Doctor) and one which does not have access to external commonsense resources (GPT).
@@
is a baseline model representing the capability of GPT-3.5 for dialogue response generation without any direct integration of commonsense. The prompt is shown in Table 7 (Appx. A).
For fair comparison, we employ the same GPT-3.5 version in each approach, choosing the latest version at the time of writing (gpt-3.5-turbo-0125) with a temperature of $0.7$.
@@ 4.1 Test Data
To conduct a realistic evaluation of the approaches under study, the test data consists of dialogues sampled from an “out-of-distribution” dataset for all models: Reflect Zhou et al. (2022a). The Reflect dataset is composed of human-written dialogues that are based on descriptions of everyday situations. Table 8 (Appx. B) presents examples of the response outputs from each model.
@@ 4.2 Evaluation
We perform pairwise preference selections between pairs of dialogue approaches, where responses from two different dialogue approaches for the same dialogue context are shown to human judges who are instructed to identify which response better satisfies the indicated characteristic. Following the improvements in dialogue characteristics from earlier commonsense-augmented dialogue works Zhou et al. (2022a); Chae et al. (2023), we evaluate based on the following four criteria:
Natural:
Which response is more natural (human-like) as the next response in the dialogue?
Engaging:
Which response is more engaging and interesting as the next response in the dialogue?
Specific:
Which response contains content that is more specific and unique to the shown dialogue?
Quality:
Which response demonstrates better overall quality as the next response in the dialogue?
Human judges are also instructed to provide a freeform text explanation of their reasoning behind which response had the better quality. 100 dialogues are sampled from the test data for use in the evaluations. Three Amazon Mechanical Turkers perform each evaluation task and are paid \$0.15 USD per task. The annotators achieve moderate inter-annotator agreement, with Krippendorff’s alpha scores of 0.442, 0.560, 0.595, and 0.561 for naturalness, engagingness, specificity, and quality, respectively. The worker screening criteria and evaluation interface are provided in Appendix C.
@@ 4.3 Results
Table 2: Pairwise evaluation results showing the preference percentages. Winning models in each comparison are statistically significant for all characteristics, except where underlined (proportion test, $\alpha<0.01$).
Table 2 shows the preference results for the models under study. It is clear from the results that ConvoSense-E outperforms the alternative approaches in generating natural, engaging, and specific responses which are overall preferred as next responses in dialogue. The preference for ConvoSense-E is particularly pronounced regarding engagingness, specificity, and overall response quality. The preference for ConvoSense-E is less dramatic for response naturalness, especially against ConvoSense-I, although ConvoSense-E still maintains a slight advantage.
ConvoSense-I is found to be the most competitive approach to ConvoSense-E, with the highest rate of preference wins relative to the alternative approaches. This is further corroborated by direct pairwise comparisons between ConvoSense-I and each of the alternative approaches, Doctor and GPT. The results show that ConvoSense-I outperforms both Doctor and GPT, although the preference is not as strong as that for ConvoSense-E.
We also find that GPT responses are more competitive with responses from ConvoSense-E than Doctor responses, as indicated by the rate of ConvoSense-E wins being lower against GPT than against Doctor. Interestingly, this suggests that the previous commonsense-augmented dialogue model is producing responses that are worse than GPT. We discuss this further in Section 5.5.
@@ 5 Discussion
5.1 Impact of Explicit Reasoning
The results in Table 2 overwhelmingly indicate that explicit reasoning over dialogue-relevant commonsense inferences results in more appealing responses to human judges. The proposed approach to dialogue modeling that decomposes response generation into separate reasoning and integration steps is quite advantageous. In this era of leveraging powerful LLMs to perform tasks with little-to-no direct training, work like that in this study is revealing that modularized generation strategies are superceding the end-to-end modeling paradigms that have dominated in the past.
Of all of the dialogue characteristics under study, the naturalness of the follow-up response is least affected by the choice of commonsense reasoning strategy. When directly comparing explicit to implicit reasoning over the same commonsense inferences, as elucidated by the pairwise selections between ConvoSense-E and ConvoSense-I, there is not a strong difference with regard to response naturalness. This suggests that, although explicit reasoning aids in the engagingness and specificity of the response to its dialogue context, it is not as useful towards improving how natural the response is for the dialogue.
Furthermore, it can be seen that there is a large preference for responses from both ConvoSense-E and ConvoSense-I when compared to GPT responses. This result indicates that augmenting responses with commonsense improves on the native response generation of GPT, regardless of integration strategy (explicit vs. implicit).
@@ 5.4 Influential Aspects from Human Feedback
We utilize an automated aspect identification procedure to summarize the human-written explanations into sets of influential characteristics of the preferred response. This aspect identification procedure uses GPT-3.5 to output short phrases (one or two words) that represent each of the indicated characteristics mentioned in the explanation (Table 9; Appx. D). On an example set of 50 explanations with characteristics identified by a human annotator, this procedure obtains 0.82 precision and 0.83 recall of the characteristics.
We run this aspect identification procedure on the explanations collected from the evaluations. We then manually review the outputted individual characteristics and find that there is a high degree of distinct yet synonymous characteristics being outputted. To aggregate the characteristics by synonymous meaning, we construct a mapping between the outputted characteristics to 12 categories. Table 8 (Appx. B) shows examples of the predicted aspects and their corresponding mapped categories. Figure 4 shows the distribution of categories in the explanations for each winning model.
@@ 5.5 Baseline Prompt Variation
One unexpected outcome of this work is the poor performance of the Doctor approach. As noted in Section 4.2, GPT is more competitive with ConvoSense-E than Doctor in terms of human preference selection on various characteristics. This is contrary to the results presented in Chae et al. (2023), which indicate that Doctor outperforms GPT in terms of naturalness and specificity. To explicitly compare the GPT baseline used in this work against Doctor, we conduct the pairwise preference selection evaluation from Section 4.2 between these two approaches. From the results shown in Table 3, it is confirmed that Doctor is indeed outperformed by our GPT baseline.
@@
Table 3: Pairwise evaluation results showing the win percentages. All wins are statistically significant (proportion test, $\alpha<0.01$).
@@
Table 4: Pairwise evaluation results showing the win percentages. All wins are statistically significant, except where underlined (proportion test, $\alpha<0.01$).
@@ 8 Ethical Considerations
Appendix A Prompts
Tables 5, 6, and 7 show the prompts for ConvoSense-E, ConvoSense-I, and GPT. For all prompts, the dialogue context is provided as a sequence of turns, separated by newlines, where each turn is prefixed by a nominal tag. The terminal turn is prefixed as “Speaker (Other)”, and its preceding utterance is labeled as “Listener (You)”. These nominal tags are then assigned in alternating order to the remainder of the turns and serve to encourage GPT-3.5 to take on the role of the Listener in the conversation when constructing the follow-up response.
Appendix B Output Examples
Examples of the response outputs, commonsense, evaluation explanations, predicted aspects, and mapped categories are shown in Table 8 for the models under study.
Appendix C MTurk Evaluation Interface
Our pilot studies on MTurk reveal that standard worker filtering criteria (Location: USA,CAN; HIT approval $\geq$ 98%; approved HITs: $\geq$ 10000) fail to ensure good-faith workers for our evaluation tasks, often yielding invalid or nonsensical explanations. To address this, we introduce a “screening” task identical to the evaluation task, used to identify reliable workers based on their written explanations. This process identifies over thirty MTurkers who provide detailed explanations for their responses. These selected workers then perform the evaluation tasks for our dialogue approaches. The preference evaluation interface used in Section 4.2 is shown in Figure 5.
Appendix D Automated Aspect Identification

### v2

@@ 1 Introduction
As such, the goal of commonsense-augmented dialogue modeling is to leverage useful commonsense inferences for producing more compelling and human-like responses in human-computer dialogue. For a given dialogue situation, there are numerous commonsense inferences that can be drawn since inferential commonsense has a many-to-many mapping due to its speculative nature Shen et al. (2022); Finch and Choi (2024). Consequently, commonsense-augmented dialogue modeling is a reasoning process of (1) deriving commonsense knowledge that likely holds true for a given dialogue context, (2) identifying a subset of true commonsense that is appropriate for generating a response, and (3) synthesizing a response from the identified commonsense knowledge. No previous work tackles all three of these components of commonsense-augmented dialogue modeling in an explicit manner. Often, the identification of appropriate commonsense for response generation is tackled jointly with response generation itself, where a generative model’s attention mechanisms implicitly learn which commonsense knowledge is most relevant to generating a response in each dialogue context Zhang et al. (2020); Sabour et al. (2022); Liu and Kilicoglu (2023). However, recent work in a variety of text generation tasks suggests that breaking down a model into a sequence of explicit reasoning steps improves the correctness and overall quality of model outputs Wei et al. (2022); Chae et al. (2023); Lee et al. (2023).
@@
We find strong evidence that explicit reasoning over commonsense inferences leads to better dialogue interactions, improving the naturalness, engagingness, specificity, and overall quality of the responses against several alternative strategies for dialogue response generation. Further analysis provides insights on which types of commonsense are most useful for response generation, revealing that predictions of personal characteristics and likely future events lead to the most favorable outcomes. Moreover, we assess human-provided explanations of response quality to identify the key response attributes that shape human perception of response quality and the particular response traits enhanced through commonsense integration. Taken together, our results demonstrate that isolating commonsense reasoning into explicit reasoning steps improves several aspects of response quality, achieving a new state-of-the-art in commonsense-augmented dialogue modeling. We make all code, models, and data publicly available to facilitate future work. [^1] https://github.com/emorynlp/DialCS-Reasoning
@@ 2 Related Work
To explore the impact of explicit reasoning over commonsense inferences against the typically utilized implicit reasoning approach, we develop two prompt-based LLM strategies for response generation that treat generated inferences as speculative thoughts to guide follow-up response generation. These two approaches utilize the same inference generation procedure and differ on the strategy for integrating commonsense inferences into a follow-up response (Figure 2). The implicit reasoning variant provides all commonsense inferences as input to an LLM which is prompted to consider the inferences when generating the best follow-up response, similar to previous works. The explicit reasoning variant involves a three-step generate-select-and-respond procedure, using LLM prompting to explicitly identify the best commonsense inferences and subsequently synthesize them into a follow-up response. We first discuss the shared Inference Generation module (§3.1), before detailing the implicit reasoning (§3.2) and the explicit reasoning approaches (§3.3).
@@ 3.1 Inference Generation
10 commonsense inferences are outputted for a provided dialogue context, each corresponding to one of the 10 commonsense types covered by ConvoSenseGenerator (Table 1). Initially, we explored outputting the top-ranked inference from beam search for each type but found significant semantic overlap in the inferences outputted across types. Since semantically unique inferences are critical for studying the impact of reasoning over these inferences, we implement a diverse beam search approach Vijayakumar et al. (2016) to output five inferences per type and then select one inference per type such that between-type cosine similarity of SBERT embeddings Reimers and Gurevych (2019) is minimized, following Finch and Choi (2024) who show diverse beam search increases inference uniqueness.
@@ 3.2 Implicit Reasoning
Given the output from inference generation, the implicit reasoning approach immediately performs response generation by taking all generated inferences as input. Table 2 provides the prompt for this approach, referred to as ConvoSense-I.
@@ 3.3 Explicit Reasoning
Given the output from inference generation, the explicit reasoning approach selects relevant inferences before composing the follow-up response. Table 3 provides the prompts for this approach, referred to as ConvoSense-E.
@@ 3.4 Prompt Formatting
For all GPT prompts, the dialogue context is provided as a sequence of turns, prefixed with speaker labels. 10 few-shot examples are provided to both ConvoSense-E and ConvoSense-I. For ConvoSense-E, we construct 10 inference selection examples, one for each inference type included in this study, and 100 response generation examples, 10 for each inference type. During inference, the 10 inference selection examples are used in the inference selection prompt. After inference selection, then the 10 response generation examples corresponding to the selected inference type are used in the response generation prompt. For ConvoSense-I, we choose 10 response generation examples from those crafted for ConvoSense-E, ensuring one example for each commonsense type. During inference, these 10 response generation examples are used in the response generation prompt.
@@ 4 Experiments
To study the impact of explicit reasoning over commonsense inferences on response generation, we compare ConvoSense-E against three alternative approaches, two which utilize implicit reasoning over commonsense (ConvoSense-I and Doctor) and one without access to external commonsense (GPT).
@@
is a model representing the baseline capability of GPT-3.5 for dialogue response generation. No commonsense or commonsense-derived response examples are provided in order to elicit the natural tendencies of GPT. The prompt is shown in Table 4.
We employ the same GPT-3.5 version in each approach, choosing the latest version at the time (gpt-3.5-turbo-0125) and temperature of $0.7$.
@@ 4.1 Test Data
To conduct a realistic evaluation of the approaches under study, the test data consists of dialogues sampled from an “out-of-distribution” dataset for all models: Reflect Zhou et al. (2022a). The Reflect dataset is composed of human-written dialogues that are based on descriptions of everyday situations. 100 dialogues are sampled for use in the evaluations, with an average of 3.1 turns and 10.8 words per utterance.
@@ 4.2 Evaluation
Figure 3: Evaluation interface.
To capture differences in end-user impressions of the different dialogue approaches, we perform pairwise preference selections between pairs of dialogue approaches using human judges. Responses from two different dialogue approaches for the same dialogue context are shown to human judges who are instructed to identify which response better satisfies the indicated characteristic. Following the improvements in dialogue characteristics from earlier commonsense-augmented dialogue works Zhou et al. (2022a); Chae et al. (2023), we evaluate the response characteristics shown in Figure 3: Naturalness (Q1), Engagingness (Q2), Specificity (Q3), and Overall Quality (Q4). Human judges are also instructed to provide a freeform text explanation of their reasoning behind which response had the better quality.
Three Amazon Mechanical Turkers perform each evaluation task and are paid \$0.15 USD per task. Our pilot studies on MTurk reveal that standard worker filtering criteria (Location: USA,CAN; HIT approval $\geq$ 98%; approved HITs: $\geq$ 10000) fail to ensure good-faith workers for our evaluation tasks, often yielding invalid or nonsensical explanations. To address this, we introduce a “screening” task identical to the evaluation task, used to identify reliable workers based on their written explanations. This process identifies over thirty MTurkers who provide detailed explanations for their responses. These selected workers then perform the evaluation tasks for our dialogue approaches and achieve moderate inter-annotator agreement, with Krippendorff’s alpha scores of 0.442, 0.560, 0.595, and 0.561 for naturalness, engagingness, specificity, and quality, respectively.
DC Speaker (Other): I had to kick Jan out of my house last night. Listener (You): What got you so mad that you kicked her out of the house? Speaker (Other): She kept bugging the dog and bothering him.
Doctor That’s not a good reason to kick someone out.
CS-I That must have been a tough situation. I hope the dog is okay. Do you think it’s time for Jan to find a new place to stay?
CS-E I can understand why you’d be upset if Jan was bothering your dog. Your home should be a peaceful and safe space for everyone, including your furry friend. How is the dog doing now that Jan’s not around?
Table 5: Example response outputs for a dialogue context (DC) from each model. CS-E/I denote ConvoSense-E/I.
@@ 4.3 Results
Table 6: Pairwise evaluation results showing the preference percentages. Winning models in each comparison are statistically significant for all characteristics, except where underlined (proportion test, $p<0.01$).
Table 5 presents examples of the response outputs from each model [^2] Full examples including commonsense generations can be found in Appx. A. and Table 6 shows the preference results for the models based on the human evaluation. These evaluation results overwhelmingly indicate that explicit reasoning over dialogue-relevant commonsense inferences results in more appealing responses to human judges. This approach consistently generates responses that are preferred by human judges for their engagingness, specificity, and overall quality, although the naturalness of the follow-up response is least affected by the choice of commonsense reasoning strategy. When directly comparing explicit to implicit reasoning over the same commonsense inferences, as elucidated by the pairwise selections between ConvoSense-E and ConvoSense-I, there is not a strong difference with regard to response naturalness. This suggests that, although explicit reasoning aids in the engagingness and specificity of the response to its dialogue context, it is not as useful towards improving how natural the response is for the dialogue. Overall, in this era of leveraging powerful LLMs to perform tasks with little-to-no direct training, work like that in this study is revealing that modularized generation strategies can outperform end-to-end modeling paradigms that have dominated in the past.
Furthermore, ConvoSense-I is found to be the most competitive approach to ConvoSense-E, with the highest rate of preference wins relative to the alternative approaches. This is further corroborated by direct pairwise comparisons between ConvoSense-I and each of the alternative approaches, Doctor and GPT, although the preference is not as strong as that for ConvoSense-E. Surprisingly, we find that the Doctor responses are the least preferred, even failing to outperform GPT itself, and we explore this further in Section 5.4. Regardless, the strong performance of both ConvoSense-E and ConvoSense-I indicate that augmenting responses with appropriate commonsense improves on the native response generation of GPT.
@@ 5.3 Influential Aspects from Human Feedback
We utilize an automated aspect identification procedure to summarize the human-written explanations into sets of influential characteristics of the preferred response. This aspect identification procedure uses GPT-3.5 to output short phrases (one or two words) that represent each of the indicated characteristics mentioned in the explanation (Table 9; Appx. B). On an example set of 50 explanations with characteristics identified by a human annotator, this procedure obtains 0.82 precision and 0.83 recall of the characteristics.
We run this aspect identification procedure on the explanations collected from the evaluations. We then manually review the outputted individual characteristics and find that there is a high degree of distinct yet synonymous characteristics being outputted. To aggregate the characteristics by synonymous meaning, we construct a mapping between the outputted characteristics to 12 categories. Table 10 (Appx. A) shows examples of the predicted aspects and their corresponding mapped categories. Figure 5 shows the distribution of categories in the explanations for each winning model.
@@ 5.4 Baseline Prompt Variation
One unexpected outcome of this work is the poor performance of the Doctor approach. As noted in Section 4.2, Doctor is the least competitive approach against ConvoSense-E across all of the characteristics. This is contrary to the results presented in Chae et al. (2023), which indicate that Doctor outperforms GPT in terms of naturalness and specificity. To explicitly compare the GPT baseline used in our work against Doctor, we conduct the pairwise preference selection evaluation between them. From the results in Table 7, it is shown that Doctor is indeed outperformed by our GPT baseline.
@@
Table 7: Pairwise evaluation results showing the win percentages. All wins are statistically significant (proportion test, $p<0.01$).
@@
Table 8: Pairwise evaluation results showing the win percentages. All wins are statistically significant, except where underlined (proportion test, $p<0.01$).
@@ 8 Ethical Considerations
Appendix A Output Examples
Full examples of the response outputs, commonsense, evaluation explanations, predicted aspects, and mapped categories are shown in Table 10 for the models under study.
Appendix B Automated Aspect Identification